OpenStudy (anonymous):

what is the slope, y-intercept, and x-intercept of -10x-7y+2=0

OpenStudy (anonymous):

translate as "solve \(-10x-7y+2=0\) for \(y\) "

OpenStudy (anonymous):

you know how to do that or you need help?

OpenStudy (anonymous):

I need help

OpenStudy (anonymous):

m= -10/7,b =2/7

OpenStudy (anonymous):

\[-10x-7y+2=0\] add \(10x\) \[-7y+2=10x\] subract \(2\) \[-7y=10x-2\] divide by \(-7\) \[y=-\frac{10}{7}x+\frac{2}{7}\]

OpenStudy (anonymous):

making \[m=-\frac{10}{7}, b=\frac{2}{7}\]

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

I put that in and it is saying that it is incorrect

OpenStudy (anwaarullah):

using the defination of intercepts when x intercpets ...x=0 then y=2/7 and similarly for y intercepts,, x=0 x=1/5 ,,,is it

OpenStudy (triciaal):

@heavenfun13 please stop giving answers without helping the person asking the question.